OpenStudy (torialaw192):

Mike observed that 75% of the students of a school liked skating. If 35 students of the school did not like skating, the number of students who liked skating was ______. (only put numeric values, no other symbols)

OpenStudy (triciaal):

total is always 100%. 2 groups those who like and those who don't like. 75% like therefore those who don't like = 100% - 75% = 25% told don't like = 35 students therefore 25% of total students, S =35 solve for S like skating = 75% of S same as S - 35

OpenStudy (torialaw192):

Ok.

OpenStudy (torialaw192):

Will give a medal!!!

OpenStudy (triciaal):

you are welcome always remember 100% is the whole/ total

OpenStudy (arnavguddu):

75 % liked skating. so (100-75) % does not like skating and that is35 so 25 % = 35 find 100 % = ????? = total number of students find 75 % = ???? = number of students who like skating

OpenStudy (arnavguddu):

also not 25 % = 35 so 1 % = 35/25 100 % = (35/25) * 100 = ???? = total 75 % = (35/25) * 75 = ???? = liked skating